# Thinking Outside the (Gray) Box: A Context-Based Score for Assessing Value and Originality in Neural Text Generation

### v1

@@ 5 Implementation and Optimization with Autoregressive Models
Calculating $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$ is not trivial. Since LLMs are trained to complete text sequences, it is unlikely that they would generate the source text immediately after the target text (which, we should remember, is generated immediately after the source text). To address this, we consider an approximation $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y}^{\prime})$, where $\mathbf{y}^{\prime}=\mathbf{y}+\mathbf{q}$. Here, $\mathbf{q}$ represents an additional question, such as “How would you describe this text?” or a similar formulation designed solely to increase the likelihood of generating the source text $\mathbf{x}$ (as well as alternative sources). Once the CoVO score has been defined, its adoption in an RL framework is straightforward. As introduced in Section 3.2, we can employ our CoVO score as the final reward for the generated sequence. Then, the model can be trained with any policy gradient method. Our experiments leverage PPO Schulman et al. (2017), which is a popular choice for training language models.
Finally, we propose an alternative method to optimize agents based on our CoVO score. While the normalization schemes mentioned earlier should help balance the value and originality components, finding the ideal reward formulation that effectively addresses both aspects is challenging. In addition, RL tends to have slower convergence. For these reasons, we propose to use DPO Rafailov et al. (2023) as an alternative solution. An intuitive implementation would consider the CoVO score from Equation 11 to rank the generated outputs for a given input. However, as already discussed, the CoVO score is made of two distinct parts that might be problematic to optimize at the same time by simply summing them up. Therefore, we propose to build two distinct rankings, one on the value component, and one on the originality component. Then, the two rankings are merged together, and the best-ranked output is the chosen one (i.e., the one whose probability will be maximized), while the worst-ranked is the rejected one (i.e., the one whose probability will be minimized). Once a batch of chosen-rejected pairs has been collected, the model can be optimized via Equation 2. Algorithm 1 summarizes the proposed training process.
Algorithm 1 DPO train step toward CoVO Score
Require $p_{\boldsymbol{\theta}}$ reference LM, $p_{\boldsymbol{\theta}^{\prime}}$ LM to be trained, =$\mathcal{B}=\{\mathbf{x}_{b}\}_{b=1}^{B}$ batch of training prompts, $K$ candidates to generate, $\mathtt{get\_ext\_reward}$ function to include potential task-specific rewards.
for $b=1\ldots B$ do
     for $k=1\ldots K$ do
         $\mathbf{y}_{k}\sim p_{\boldsymbol{\theta}^{\prime}}(\mathbf{x}_{b})$
         $\mathsf{value}_{k}=s_{v}(\mathbf{x}_{b},\mathbf{y}_{k},p_{\boldsymbol{\theta}})$
         $\mathsf{orig}_{k}=s_{o}(\mathbf{x}_{b},\mathbf{y}_{k},p_{\boldsymbol{\theta}})$
         $\mathsf{ext\_reward}_{k}=\mathtt{get\_ext\_reward}(\mathbf{y}_{k})$
     end for
     $\mathsf{new\_scores}=0\ldots K-1$
     _$\_,\mathsf{ids}=\mathtt{sort}(\mathsf{value})$
     $\mathsf{value}=\mathtt{sort\_by\_idx}(\mathsf{new\_scores},\mathsf{ids})$
     _$\_,\mathsf{ids}=\mathtt{sort}(\mathsf{orig})$
     $\mathsf{orig}=\mathtt{sort\_by\_idx}(\mathsf{new\_scores},\mathsf{ids})$
     $\mathsf{score}=\mathsf{value}+\mathsf{orig}$
     $\mathsf{score}=\mathsf{score}+\mathsf{ext\_reward}$
     $\mathsf{chosen}_{b}=\mathbf{y_{\mathtt{argmax}(\mathsf{score})}}$
     $\mathsf{rejected}_{b}=\mathbf{y_{\mathtt{argmin}(\mathsf{score})}}$
end for
Train $p_{\boldsymbol{\theta}^{\prime}}$ via Equation 2 on $\mathsf{chosen}$ and $\mathsf{rejected}$.
6 Experiments
@@
We evaluate the effectiveness of our RL strategy through two case studies: poetry generation (Section 6.1) and mathematical problem resolution (Section 6.2) [^1] The code for the experiments will be made available with the camera-ready version of the paper if accepted..
6.1 Poetry Generation
@@
Llama3- Correctness $\uparrow$ Metric (L/S) $\uparrow$ T-LCS (avg/max) $\downarrow$ Correctness $\uparrow$ Metric (L/S) $\uparrow$ T-LCS (avg/max) $\downarrow$
Baseline $\mathbf{1.00}$ $\mathbf{0.60}$ / $\mathbf{0.60}$ $8.0$ / $\underline{57}$ $\mathbf{1.00}$ $0.33$ / $\underline{0.30}$ $5.0$ / 88880
CoVO $\mathbf{1.00}$ $\mathbf{0.60}$ / $\mathbf{0.60}$ $\underline{9.9}$ / $49$ $\mathbf{1.00}$ $\mathbf{0.53}$ / $\underline{0.30}$ $7.2$ / $\underline{41}$
CoVO-std $\underline{0.76}$ $\underline{0.20}$ / $\underline{0.33}$ $5.8$ / $19$ $\underline{0.40}$ $\underline{0.27}$ / $\mathbf{0.57}$ $\mathbf{4.8}$ / $\mathbf{6}$0
CoVO-run $0.80$ $0.40$ / $0.57$ $\mathbf{4.9}$ / $\mathbf{7}$0 $0.60$ $0.47$ / $0.40$ $\mathbf{4.8}$ / 77770
CoVO-dpo $0.96$ $0.50$ / $0.50$ $6.1$ / $14$ $0.92$ $0.47$ / $0.40$ $\underline{7.8}$ / $40$
@@
Table 1: Aggregate results of CoVO scores at inference time considering both training prompts (left) and testing prompts (right). Scores on the poetical metrics are reported at the line level (L) and syllable level (S) and only consider requests for styles with specific metrical properties. The best scores are highlighted in bold, while the worst scores are indicated with underlining.
6.1.1 Experimental Setup
The first set of experiments aims to teach the LLM to generate poems that are both more original and valuable. More specifically, we follow the approach outlined by Bradley et al. (2024) and instruct the model to write a poem in a particular style and tone. We consider the Llama3-8B model Grattafiori et al. (2024) as our pre-trained agent. Since we do not use the instruction-tuned model, we prompt it with some few-shot examples of the task to make it more likely to produce the desired output in the desired form (see the full prompt in Appendix A). Instead of fine-tuning the entire network, we consider Low-Rank Adaptation (LoRA) Hu et al. (2022). The original model is also used to compute the score. We experiment with various settings, i.e., with Equation 11 and $\lambda_{v}=\lambda_{o}=1.0$ (hereinafter Llama3-CoVO); by standardizing the two score portions separately at the batch level (Llama3-CoVO-std); by standardizing them using running statistics (Llama3-CoVO-run); and by using the DPO adaptation through Algorithm 1 (Llama3-CoVO-dpo). Due to additional resource consumption, the DPO adaptation considers a 4-bit quantization of the model Dettmers and Zettlemoyer (2023). The full training parameters are reported in Appendix A. We consider the pre-trained version as our baseline, and we perform a quantitative evaluation. In particular, we compute poetical metrics for quality (lexical correctness of poems and adherence to line- and syllable-level constraints) and for originality (accidental reproduction of existing poems). For the latter, we define a Token-based Longest Common Substring (T-LCS) score, and we use it by comparing generated poems with a reference dataset of approx. 84k public-domain poems extracted from Project Gutenberg (see Appendix B for a first presentation of our GutenVerse dataset). While a generated poem can be an accidental reproduction of a protected work or a different kind of text (e.g., a song), we believe it can provide a useful evaluation tool to understand the general degree of originality.
6.1.2 Experimental Results
@@
Table 1 reports the scores about the compliance of poetical constraints at the syllable and line levels, lexical correctness (as the ratio of poems not containing noisy text), and accidental reproduction rate (as the mean and maximum token-based longest common substring).
As expected, different training strategies have different effects on the aforementioned metrics. In particular, using the CoVO score as-is leads to behavior close to the baseline for training tone-style pairs, while it increases metric adherence but also verbatim reproduction for out-of-distribution pairs. Llama3-CoVO-dpo has similar but less extreme results, without excelling or failing in any metric. On the other hand, standardizing the CoVO score components places more importance on exploration: both Llama3-CoVO-std and Llama3-CoVO-run tend to produce more noisy outputs but without any relevant accidental reproduction.
@@
Interestingly, these considerations align well with our CoVO score. Figure 1 reports the value and originality according to Equation 11 under the pre-trained model. While the different methods do not significantly differ from the baseline (which is possibly due to the opposite forces of value and originality Varshney (2019)), we again see that using the CoVO score as-is or with DPO places more focus on the value part, without substantial increases in originality. Instead, standardizing the two separate parts either at the batch or the training level reverses the situation, with better performances on originality but slightly worse on value. However, aggregated scores, such as those presented here, might be insufficient. For a more complete overview, we also conducted a more fine-grained analysis of the generated poems in Appendix C.
@@
6.2 Math Problem Resolution
6.2.1 Experimental Setup
@@
The second set of experiments aims to teach the LLM to solve mathematical problems through more diverse procedures. In particular, we focus on the Mistral-based Jiang et al. (2023) MetaMath model, i.e., fine-tuned with self-supervised learning on the MetaMathQA Yu et al. (2024). It is a dataset of textual math questions paired with responses where the numerical answer is easily separable from the textual procedure. While the entire set contains 395k entries, making an additional training epoch too expensive, MetaMathQA is composed of entries from two different training sets, then augmented with various techniques: GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021). Since we are only interested in the questions, we limit our training to those datasets. Moreover, we exclude all questions with a tokenized length of either question or answer greater than 512, obtaining 14876 out of 14973 total entries.
@@
We separate the procedure and the answer from each solution to train our model and use the numerical answer to check the correctness of the predicted solution. Because of this, the RL problem can be formulated considering up to two rewards: our CoVO score computed on the procedure and an extrinsic reward based on the correctness of the answer. As for the previous case study, instead of fine-tuning the entire model, we adopt a more parameter-efficient strategy with LoRA and use the original model to perform the CoVO score computation. Again, we experiment with four different configurations: PPO and the score from Equation 11 with $\lambda_{v}=\lambda_{o}=1.0$ (MetaMath-mistral-CoVO); PPO and the score normalized at the batch level (MetaMath-mistral-CoVO-std); PPO and the score normalized at training level (MetaMath-mistral-CoVO-run); and the DPO strategy (MetaMath-mistral-CoVO-dpo). We consider scenarios with and without the external reward, and we compare the performances with the original model and a fine-tuning based only on the external reward. For reproducibility, the full training parameters are reported in Appendix A.
@@
The evaluation considers both GSM8K and MATH test sets (limited to the entries with a tokenized length of question and answer smaller than 512, i.e., all 1319 entries for GSM8K and 4546 out of 5000 for MATH). We compute the percentage of correct solutions together with two diversity metrics: expectation-adjusted distinct N-grams (EAD) Liu et al. (2022) and sentence embedding cosine similarity (Sent-BERT) Hong et al. (2024), which should measure syntactical and semantical diversity, respectively Kirk et al. (2024). EAD counts the number of distinct N-grams (averaging over $N=1\ldots 5$) across all generated responses and removes the bias toward shorter inputs by scaling the number of distinct tokens based on their expectations. The Sent-BERT metric computes the average of the cosine similarity between the embeddings of any possible pairs of outputs and returns 1 minus the similarity. This was originally based on Sentence-BERT Reimers and Gurevych (2019), we employ instead the more recent all-mpnet-base-v2, as suggested by their developers [^2] https://huggingface.co/sentence-transformers/bert-large-nli-stsb-mean-tokens.
@@
Following Kirk et al. (2024), we compute cross-input EAD and Sent-BERT, i.e., we derive them by considering all outputs produced for a specific seed together. In addition, we also calculate against-pretrained EAD and Sent-BERT. Given each input, we compare the output with the one from the pre-trained model by calculating the average expectation-adjusted distinct N-grams not present in the pre-trained model response, and 1 minus the cosine similarity between the two outputs, respectively.
@@
MetaMath-mistral placeholder EAD $\uparrow$ Sent-BERT $\uparrow$ EAD $\uparrow$ Sent-BERT $\uparrow$
Baseline $77.96\%(3)$ $2.007$ $\underline{0.732}$
Ext. reward $78.32\%(0)$ $1.995$ $0.738$ $0.042\pm.005$ $0.109\pm.011$
CoVO $\mathbf{78.97\%}(2)$ $1.995$ $0.735$ $0.047\pm.005$ $0.125\pm.011$
CoVO w ext $78.68\%(1)$ $2.000$ $0.736$ $0.039\pm.005$ $0.107\pm.010$
CoVO-std $78.89\%(2)$ $1.992$ $0.736$ $0.041\pm.004$ $0.115\pm.011$
CoVO-std w ext $78.47\%(0)$ $1.985$ $\mathbf{0.740}$ $0.039\pm.004$ $0.106\pm.010$
CoVO-run $78.42\%(3)$ $1.993$ $0.734$ $0.044\pm.005$ $0.119\pm.011$
CoVO-run w ext $78.17\%(0)$ $1.988$ $0.739$ $0.043\pm.005$ $0.114\pm.011$
CoVO-dpo $77.85\%(5)$ $2.017$ $0.736$ $\underline{0.028\pm.004}$ $\underline{0.075\pm.009}$
CoVO-dpo w ext $\underline{77.63\%}(5)$ $\mathbf{2.019}$ $0.733$ $\underline{0.025\pm.004}$ $\underline{0.066\pm.009}$
@@
MetaMath-mistral placeholder EAD $\uparrow$ Sent-BERT $\uparrow$ EAD $\uparrow$ Sent-BERT $\uparrow$
Baseline $33.55\%(483)$ $5.724$ $\underline{0.654}$
Ext. reward $\underline{32.82\%}(411)$ $5.666$ $0.658$ $0.118\pm.004$ $0.176\pm.007$
CoVO $33.38\%(424)$ $\underline{5.643}$ $0.663$ $\mathbf{0.134\pm.005}$ $\mathbf{0.200\pm.007}$
CoVO w ext $32.92\%(397)$ $5.647$ $0.661$ $0.123\pm.005$ $0.184\pm.007$
CoVO-std $33.24\%(433)$ $5.676$ $0.660$ $\mathbf{0.128\pm.005}$ $\mathbf{0.190\pm.007}$
CoVO-std w ext $33.38\%(409)$ $5.646$ $\underline{0.654}$ $0.123\pm.004$ $\mathbf{0.187\pm.007}$
CoVO-run $32.91\%(413)$ $5.671$ $\mathbf{0.665}$ $\mathbf{0.134\pm.005}$ $\mathbf{0.199\pm.007}$
CoVO-run w ext $33.24\%(388)$ $5.654$ $\underline{0.654}$ $\mathbf{0.129\pm.005}$ $\mathbf{0.195\pm.007}$
CoVO-dpo $33.62\%(474)$ $\mathbf{5.753}$ $\underline{0.654}$ $0.072\pm.004$ $0.105\pm.006$
CoVO-dpo w ext $\mathbf{34.00\%}(493)$ $5.747$ $0.655$ $\underline{0.063\pm.004}$ $\underline{0.093\pm.005}$
@@
6.2.2 Experimental Results
Tables 2 and 3 report the results for the GSM8K and MATH test sets. For the GSM8K test set, while all tested methods achieve similar results, using the extrinsic reward leads to higher Sent-BERT scores but lower EAD and solved problems. In particular, combining it with DPO yields the lowest accuracy and the smallest deviations from the original model, whereas the ‘plain’ CoVO reward leads to the highest percentage of correct answers.
The results for the MATH test set are significantly different. The CoVO-dpo strategy with extrinsic rewards achieves this time the highest percentage of solved problems, despite the extrinsic reward alone being the worst solution. The DPO strategy tends also to enhance EAD diversity, but causes little deviation from the pre-trained model; in contrast, the other three methods improve Sent-BERT diversity but obtain performances worse than the baseline model. Notably, the extrinsic reward leads to fewer variations from the original model, while not providing significant improvements in accuracy. However, the number of unfinished responses is quite high (approx. 1 out of 10), and results might significantly vary if more tokens are allowed during generation.
7 Conclusion
In this paper, we have presented CoVo, a novel score that quantifies the value and originality of neurally-generated text. The definition of CoVO is based on the analysis of the mutual information between the model’s outputs and inputs, and vice versa. We have also proposed an optimization problem where a generative model aims to maximize this score in order to generate more creative products, and we have defined several strategies to use it in the context of language modeling. We have conducted experiments on poetry generation and math problem resolution, exploring the trade-offs in terms of accuracy vs diversity.
Effectively balancing value and originality maximization remains an open question, but our score seems to appropriately correlate with domain-specific measures. Our research agenda aims to extend this scoring system to other models and tasks and to explore its use for evaluation rather than solely for optimization. We also plan to investigate the definition of additional scores for capturing other potentially relevant aspects of the creative process.
8 Limitations

### v2

@@ A Context-Based Score for Valuable and Original Generation
Calculating $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$ is not trivial. Since LLMs are trained to complete text sequences, it is unlikely that they would generate the source text immediately after the target text (which, we should remember, is generated immediately after the source text). To address this, we consider an approximation $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y}^{\prime})$, where $\mathbf{y}^{\prime}=\mathbf{y}+\mathbf{q}$. Here, $\mathbf{q}$ represents an additional question, such as “How would you describe this text?” or a similar formulation designed solely to increase the likelihood of generating the source text $\mathbf{x}$ (as well as alternative sources). Once the CoVO score has been defined, its adoption in an RL framework is straightforward. As previously introduced, we can directly utilize our CoVO score as the final reward for the generated sequence.
Then, the model can be trained with any policy gradient method. Our experiments leverage GRPO (Shao et al. 2024), which is a state-of-the-art choice for training language models. As introduced above, GRPO adds a per-token KL divergence term to the loss rather than to the single rewards. Usually, the KL divergence is approximated with the following estimator (Schulman 2020): $r_{ref}(\boldsymbol{\theta})-\log r_{ref}(\boldsymbol{\theta})-1$, where $r_{ref}(\boldsymbol{\theta})=\frac{\pi_{ref}}{\pi_{\boldsymbol{\theta}}}$ and $\pi_{ref}$ is the reference policy, i.e., the model before GRPO training. In particular, GRPO aims to minimize the KL divergence, thus the second term $-\log r_{ref}=-\log\pi_{ref}+\log\pi_{\boldsymbol{\theta}}$ can be seen as made of two optimization problems: the minimization of the originality component $-\log\pi_{ref}$ from Equation 9, and the maximization of surprisal, or self-information, $-\log\pi_{\boldsymbol{\theta}}$ under the current model. In other words, this second term somehow trades off a portion of the originality component under the reference model (proportional to the $\beta\ll 1$ coefficient) with the surprisal under the current model. However, the first term of the KL approximation keeps the two policies closer, preventing the trained one from deviating too much and potentially disrupting the impact of the originality component in favor of the value component.
Experiments
@@
We evaluate the effectiveness of our RL strategy through three case studies: poetry generation, mathematical problem resolution, and the tasks included in NoveltyBench [^1] The code and results of the experiments can be found at
https://anonymous.4open.science/r/CoVO-grpo/
. In all experiments, we employ two settings, i.e., GRPO to maximize the score from Equation 11 without the KL divergence loss, i.e., with $\beta=0.0$ (CoVO); and GRPO to maximize the score from Equation 11 with the KL divergence loss, i.e., with $\beta=0.05$ (CoVO + KL). Both methods assume $\lambda_{v}=\lambda_{o}=1.0$.
Poetry Generation
@@
Corr. $\uparrow$ Metric (L/S) $\uparrow$ T-LCS $\downarrow$ Tone $\uparrow$ Corr. $\uparrow$ Metric (L/S) $\uparrow$ T-LCS $\downarrow$ Tone $\uparrow$
Meta-Llama-3-8B 0.9870.9870.987 0.3000.3000.300 / 0.1770.1770.177 6.0676.0676.067 / 191919 $0.664_{\pm 0.069}$ 1.0001.0001.000 0.4440.4440.444 / 0.1320.1320.132 5.8535.8535.853 / 686868 $0.621_{\pm 0.063}$
+ CoVO 1.0001.0001.000 0.2000.2000.200 / 0.0840.0840.084 4.8804.8804.880 / 0777 $0.749_{\pm 0.055}$ 0.9600.9600.960 0.2670.2670.267 / 0.0460.0460.046 4.9334.9334.933 / 0666 $0.661_{\pm 0.060}$
+ CoVO + KL 0.9870.9870.987 0.3330.3330.333 / 0.0660.0660.066 5.1075.1075.107 / 0999 $0.745_{\pm 0.057}$ 0.9730.9730.973 0.2890.2890.289 / 0.1170.1170.117 5.1605.1605.160 / 222222 $0.682_{\pm 0.059}$
@@
Table 1: Aggregate results of CoVO scores at inference time considering both training prompts (left) and testing prompts (right). Scores on the poetical metrics are reported at the line level (L) and syllable level (S) and only consider requests for styles with specific metrical properties. Under T-LCS, we report both the mean and the maximum longest common substring across all generated poems. The mean and the 95% confidence interval are reported for tone adherence.
Experimental Setup.
[…]
Experimental Results.
@@
Table 1 reports the scores about the compliance of poetical constraints at the syllable and line levels, lexical correctness (as the ratio of poems not containing noisy text), tone adherence (as the zero-shot classification of that poem being of that tone rather than its opposite), and accidental reproduction rate (as the mean and maximum token-based longest common substring).
Overall, our CoVO-based fine-tuning leads to a higher tone adherence and lower reproduction rate, at the potential cost of metric adherence, especially without the KL loss. Indeed, its role seems to foster quality (especially in terms of metrical correctness), trading off some originality. On the contrary, not using the KL loss arguably avoids any significant reproduction, as demonstrated by the very low maximum token-based longest common substring.
@@
Interestingly, these considerations align well with our CoVO score. Figure 2 reports the value and originality according to Equation 11 under the pre-trained model. While the two methods do not significantly differ from the baseline (which is possibly due to the opposite forces of value and originality (Varshney 2019)), we again see that the presence of KL leads to slightly higher value, while its absence leads to slightly higher originality. However, aggregated scores, such as those presented here, might be insufficient. For a more complete overview, we also conducted a fine-grained analysis of the generated poems in the Technical Appendix.
@@
Math Problem Resolution
Experimental Setup.
@@
The second set of experiments concerns a more practical and quantitative task, as it aims to teach the LLM to solve mathematical problems through more diverse procedures. In particular, we focus on the Mistral-based (Jiang et al. 2023) MetaMath-Mistral-7B model, i.e., fine-tuned with self-supervised learning on the MetaMathQA (Yu et al. 2024). It is a dataset of textual math questions paired with responses where the numerical answer is easily separable from the textual procedure. While the entire set contains 395k entries, making an additional training epoch too expensive, MetaMathQA is composed of entries from two different training sets, then augmented with various techniques: GSM8K (Cobbe et al. 2021) and MATH (Hendrycks et al. 2021). Since we are only interested in the questions, we limit our training to a single epoch over those datasets. Moreover, we exclude all questions with a tokenized length of either question or answer greater than 512, obtaining 14876 out of 14973 total entries.
@@
We separate the procedure and the answer from each solution to train our model and use the numerical answer to check the correctness of the predicted solution. Because of this, the RL problem can be formulated considering up to two rewards: our CoVO score computed on the procedure and an extrinsic reward based on the correctness of the answer. As for the previous case study, instead of fine-tuning the entire model, we adopt a more parameter-efficient strategy with LoRA and use the original model to perform the CoVO score computation. Following (Yu et al. 2024), the outputs are obtained with a greedy strategy.
@@
The evaluation considers both GSM8K and MATH test sets (limited to the entries with a tokenized length of question and answer smaller than 512, i.e., all 1319 entries for GSM8K and 4546 out of 5000 for MATH). We compute the percentage of correct solutions together with two diversity metrics: expectation-adjusted distinct N-grams (EAD) (Liu et al. 2022) and sentence embedding cosine similarity (SBERT) (Hong et al. 2024), which should measure syntactical and semantical diversity, respectively (Kirk et al. 2024). EAD counts the number of distinct N-grams (averaging over $N=1\ldots 5$) across all generated responses and removes the bias toward shorter inputs by scaling the number of distinct tokens based on their expectations. The SBERT metric computes the average of the cosine similarity between the embeddings of any possible pairs of outputs and returns 1 minus the similarity. This was originally based on Sentence-BERT (Reimers and Gurevych 2019), we employ instead the more recent all-mpnet-base-v2, as suggested by their developers [^2] https://huggingface.co/sentence-transformers/bert-large-nli-stsb-mean-tokens.
@@
Following Kirk et al. (2024), we compute cross-input EAD and SBERT, i.e., we derive them by considering all outputs produced for a specific seed together. In addition, we also calculate against-pretrained EAD and SBERT. Given each input, we compare the output with the one from the pre-trained model by calculating the average expectation-adjusted distinct N-grams not present in the pre-trained model response, and 1 minus the cosine similarity between the two outputs, respectively.
@@
EAD $\uparrow$ SBERT $\uparrow$ EAD $\uparrow$ SBERT $\uparrow$
[…]
+ CoVO + Ext $\mathbf{78.33}\%(4)$ 2.03402.03402.0340 $\mathbf{0.6404}$ $0.0628_{\pm.0056}$ $0.0088_{\pm.0011}$
+ CoVO + Ext + KL $\underline{77.95}\%(4)$ 2.03672.03672.0367 0.64020.64020.6402 $0.0638_{\pm.0057}$ $0.0089_{\pm.0011}$
@@
EAD $\uparrow$ SBERT $\uparrow$ EAD $\uparrow$ SBERT $\uparrow$
MetaMath-Mistral-7B 33.55%(483)33.$5$ ( 483 ) 5.72395.72395.7239 0.80320.80320.8032 - -
+ Ext $\underline{33.19}\%(469)$ $\underline{5.7187}$ $\underline{0.8027}$ $\underline{0.0333_{\pm.0029}}$ $\underline{0.0074_{\pm.0008}}$
[…]
+ CoVO + Ext $\mathbf{33.76}\%(503)$ 5.81145.81145.8114 0.80300.80300.8030 $0.1136_{\pm.0047}$ $0.0259_{\pm.0015}$
[…]
@@
Experimental Results.
Tables 2 and 3 report the results for the GSM8K and MATH test sets. For the GSM8K test set, while all methods achieve similar results, using the CoVO score only (with and without the KL loss) leads to greater EAD diversity and to diverge more from the original model, while the presence of the math reward leads to greater accuracy, especially without KL.
The results for the MATH test confirm that the most accurate method is the one trained to optimize the CoVO score and the extrinsic reward, with a negligible trade-off in terms of diversity, since the cross-input EAD and the against-pretrained scores are still substantially higher than those from the baselines. However, removing the extrinsic reward likely pushes the model too far from its pre-trained version, causing the accuracy to decrease.
Figure 3: The mean utility (quality) and mean distinct (novelty) scores on NoveltyBench (‘curated’ partition) for the original model and our methods (tuned on the ‘wildchat’ partition).
NoveltyBench
Experimental Setup.
Finally, we also experiment with NoveltyBench (Zhang et al. 2025), a very recent benchmark that aims to evaluate the ability of language models to produce multiple distinct and high-quality outputs. NoveltyBench contains two sets of prompts thought for eliciting diverse responses: the ‘curated’ partition, with 100 prompts manually curated by the paper’s authors, and the ‘wildchat’ partition, with 1000 prompts sourced from the WildChat-1M dataset (Zhao et al. 2024). The benchmark requires a language model to generate 10 outputs per prompt using a temperature setting of 1.01.01.0, after which it computes scores for novelty and quality. To assess novelty, the 10 outputs are grouped into equivalence classes using a fine-tuned DeBERTa model (He et al. 2021), and the number of distinct classes is reported as the novelty score. For the latter, it computes the cumulative utility of the 10 outputs, where the utility is zero if the $i$-th output has the same equivalence class as a precedent output, and a calibrated reward from Skywork-Reward-Gemma-2-27B-v0.2 (Liu et al. 2024) otherwise (we refer to Zhang et al. (2025) for full details). To evaluate our approach with NoveltyBench, we fine-tune the Llama-3.2-3B-Instruct model (Grattafiori et al. 2024) for a single epoch on the ‘wildchat’ partition (the full parameters are reported in the Technical Appendix); then, we compute the novelty and quality scores on the ‘curated’ partition and we compare them with the performance of the original model.
Experimental Results.
Figure 3 reports the quality and novelty scores achieved by our methods, compared to those from the original instructed model. Optimizing for the CoVO score results in substantial improvements in both novelty and quality metrics, with the greatest gains in novelty. Moreover, these results underscore the interplay between KL loss and our CoVO score: incorporating the KL penalty tends to improve quality, but at the cost of reduced novelty.
Discussion
The CoVO score captures properties that are functionally aligned with creativity-relevant aspects of language generation, as demonstrated through both conceptual analysis and empirical results. When used as a reward function in Group Relative Policy Optimization (GRPO), CoVO drives improvements in the novelty and quality of model outputs, even with a relatively small number of optimization steps. Although KL-divergence regularization is typically employed to constrain policy shifts and preserve alignment with the base model distribution, CoVO can contribute independently to several desirable behaviors: reducing the risk of inadvertent memorization of copyrighted material, promoting diversity in generated outputs, and mitigating undesirable inductive biases introduced during pretraining.